Exercise 18.108 Part A A 130.0-mL sample of a solution that is 2.8*10-3 M in AgNO2 is mixed with a 230.0 After the solution reaches equibrium, what concentration of Ag (aq) remains? mL sample of a solution that is 0.12 M in NaCN Express your answer using two significant figures. 1 无 Ag Submit My Answers Give Up
Answer #1

no of moles of AgNO3   = molaity * volume in L

                                       = 0.0028*0.13

                                      = 0.000364moles

no of moles of NaCN    = molarity * volume in L

                                      = 0.12*0.23   = 0.0276moles

         Ag^+(aq)    +     2CN^- ----------------->[ Ag(CN)2]^-

I        0.000364               0.0276                        0

C       -0.000364            2*-0.000364               0.000364

E          0                        0.026872                     0.000364

          Kf   = [Ag(CN)2]^-/[Ag^+][CN^-]^2

            5.6*10^18   = 0.000364/x*(0.026872)^2

           x                  = 0.000364/(5.6*10^18)(0.026872)^2    = 9*10^-20

   [Ag^+]    = 9*10^-20/0.13 = 6.92*10^-19 M